# Metric-Guided Conformal Bounds for Probabilistic Image Reconstruction

### v3

@@ Abstract
Modern deep learning reconstruction algorithms generate impressively realistic scans from sparse inputs, but can often produce significant inaccuracies. This makes it difficult to provide statistically guaranteed claims about the true state of a subject from scans reconstructed by these algorithms. In this study, we propose a framework for computing provably valid prediction bounds on claims derived from probabilistic black-box image reconstruction algorithms. The key insights behind our framework are to represent reconstructed scans with a derived clinical metric of interest, and to calibrate bounds on the ground truth metric with conformal prediction (CP) using a prior calibration dataset. These bounds convey interpretable feedback about the subject’s state, and can also be used to retrieve nearest-neighbor reconstructed scans for visual inspection. We demonstrate the utility of this framework on sparse-view computed tomography (CT) for fat mass quantification and radiotherapy planning tasks. Results show that our framework produces bounds with better semantical interpretation than conventional pixel-based bounding approaches. Furthermore, we can flag dangerous outlier reconstructions that look plausible but have statistically unlikely metric values.
Keywords:
Image Reconstruction Conformal Prediction Sparse-view CT Deep Generative Models
@@ 1 Introduction
Most existing strategies applying CP to image reconstruction operate directly on pixel values [2, 16, 29], which yield bounds with limited clinical interpretability. Other studies propose using low-dimensional representations of pixels computed using principal component analysis (PCA) [3] or latent spaces of pretrained generative models [27]. However, PCA is expensive to compute for scans larger than a small size, and high-quality generative models are not available for arbitrary medical image sets. In contrast, we propose computing CP bounds over metrics derived from the reconstructed scans, resulting in a flexible, interpretable, and computationally tractable solution.
@@
     $[s_{i,lo},s_{i,hi}]\leftarrow[Q_{\alpha_{lo}}(\hat{Y}_{i})-Y_{i},Y_{i}-Q_{1-\alpha_{hi}}(\hat{Y}_{i})]$
@@
$[q_{lo},q_{hi}]\leftarrow[Q_{1-\hat{\alpha}_{lo}}(\{s_{i,lo}\}_{i=1}^{n}),Q_{1-\hat{\alpha}_{hi}}(\{s_{i,hi}\}_{i=1}^{n})]$
@@ 2 Methods
For test subject $n+1$, we aim to derive bounds $C(\hat{Y}_{n+1})=[L(\hat{Y}_{n+1}),U(\hat{Y}_{n+1})]$ that contain the ground truth metric $Y_{n+1}$ with $(1-\alpha_{lo}-\alpha_{hi})$average certainty (known as marginal coverage [11]). If the reconstruction pipeline were perfectly calibrated, we could simply set $L(\hat{Y}_{n+1})=Q_{\alpha_{lo}}(\hat{Y}_{n+1})$ and $U(\hat{Y}_{n+1})=Q_{\alpha_{hi}}(\hat{Y}_{n+1})$. However,to handle mis-calibrations, we will use $\mathcal{D}$ to find appropriate adjustments to $Q_{\alpha_{lo}}(\hat{Y}_{n+1})$ and $Q_{\alpha_{hi}}(\hat{Y}_{n+1})$. We present how we find these adjustments in Sec. 2.1, and the use of the metric bounds to retrieve inlier, outlier, and nearest neighbor reconstructed scans in Sec. 2.2.
@@ 2.1 Quantile Adjustments Using Calibration Data
Using $\mathcal{D}$, we derive adjustments $q_{lo}$ and $q_{hi}$ to the $Q_{\alpha_{lo}}(\hat{Y}_{n+1})$ and $Q_{1-\alpha_{hi}}(\hat{Y}_{n+1})$ quantiles of the ground truth metric of test subject $n+1$ (Alg. 1). To do so, we use a probabilistic “split” CP setup proposed in previous work [26, 32]. We first calculate $s_{lo,i}$ and $s_{hi,i}$, the lower and upper non-conformity scores for each patient, capturing the differences between the lower and upper sample quantiles and ground truth metrics: $s_{i,lo}=Q_{\alpha_{lo}}(\hat{Y}_{i})-Y_{i}$ and $s_{i,hi}=Y_{i}-Q_{1-\alpha_{hi}}(\hat{Y}_{i})$. We then compute the $(1-\alpha_{lo})$-th and $(1-\alpha_{hi})$-th empirical quantiles of these scores to arrive at the adjustments: $q_{lo}=Q_{1-\hat{\alpha}_{lo}}(\{s_{i,lo}\}_{i=1}^{n})$ and $q_{hi}=Q_{1-\hat{\alpha}_{hi}}(\{s_{i,hi}\}_{i=1}^{n})$, where $\hat{\alpha}_{lo}=\lfloor\alpha_{lo}(n+1)\rfloor/n$ and $\hat{\alpha}_{hi}=\lfloor\alpha_{hi}(n+1)\rfloor/n$ are finite-sample adjusted mis-coverage rates.
@@ 2.2 Retrieving Scans Using Metric Bounds
While the metric bounds are useful in their own right to offer interpretable feedback on patient state, they can also be mapped back to the reconstructed scans $\{\hat{Y}_{n+1,j}\}_{j=1}^{n_{s}}$ to provide visual feedback to the user or clinician (Alg. 2). First, we can flag “outlier” scans, i.e., those scans with metrics outside $[L(\hat{I}_{n+1}),U(\hat{I}_{n+1})]$. Second, we can retrieve the scans with metrics nearest to $L(\hat{I}_{n+1})$ and $U(\hat{I}_{n+1})$, thereby providing visual interpretability to the metric bounds.
@@ 3 Experiments
Fat Mass Estimation. Quantifying thoracic fat mass distribution is crucial to understanding its role in cardiometabolic risk and other health outcomes [7, 31]. We computed fat volume for a 2D CT slice by multiplying the number of fat pixels with slice thickness. We identified fat pixels by thresholding values in [-150,-50] Hu. We used a de-identified CT dataset of 935 subjects from The University of Texas M.D. Anderson Cancer Center [^1] This research was conducted using an approved institutional review board protocol.. We split this dataset into 729 subjects ($\sim$300k axial slices) for training and 265 subjects (795 axial slices) for testing. We trained an unconditional 2D U-Net diffusion model on training data and used guided posterior sampling to reconstruct slices from 15 1-D observed projections [8, 14]. We used $n=596$ slices for calibration and 199 slices for testing, and generated $n_{s}=100$ reconstructions per slice using random input noise initializations. We repeated the experiment 1000 times with random calibration-testing splits.
@@
Table 1: Quantitative comparison of Metric CP to baselines. We used 1000 random $75\%$-$25\%$ calibration-testing splits, and report mean coverage and normalized interval lengths. The normalized interval length for a baseline method is its mean interval length divided by Metric CP’s mean interval length. While Metric, Pixel, and Pixel CP bounds often have smaller (“tighter”) interval lengths than Metric CP (a normalized length $<1$), they suffer from significant under-coverage, making them unusable for interpretable confidence assessments. By construction, Metric CP provides a target coverage with reasonable interval lengths.
@@ 4 Conclusion
We proposed a theoretically grounded, application-tuned framework to identify the quality of algorithmically reconstructed scans, and derive valid prediction intervals on derived metrics. Results show that representing scans with metrics, as opposed to operating in the raw pixel space, helps produce interpretable confidence intervals that capture clinically meaningful spatial correlations between pixels [19, 1, 6]. However, naively computing confidence intervals directly from these metrics without calibration is insufficient, as they do not accurately reflect the true distribution of ground truth values. The proposed framework could prove useful in flagging situations where a scan-derived metric is statistically likely to be outside of acceptable safety bounds. For example, maximum dose to the heart in RT planning should be $<$5Gy. If bounds for a subject indicate a good chance of the metric going beyond this limit, the case can be flagged for further review and intervention [21, 25, 33].
Our framework has a few assumptions and limitations to consider in practice. First, CP assumes exchangeability of data samples. Practitioners should ensure that calibration data is representative of future test data. Furthermore, the current method focuses on computing bounds on one metric, but there are often several metrics of interest in any application. A natural extension of this work would be to derive multi-metric bounds for image reconstruction. Finally, image bound retrieval is only useful if there are reconstructions in the scan set with metric values sufficiently close to the metric bounds. If, for example, the reconstruction algorithm has significant bias or the number of samples $n_{s}$ is small, this may not hold and lead to misleading visual interpretations.
Acknowledgements

### v4

@@ Abstract
Modern deep learning reconstruction algorithms generate impressively realistic scans from sparse inputs, but can often produce significant inaccuracies. This makes it difficult to provide statistically guaranteed claims about the true state of a subject from scans reconstructed by these algorithms. In this study, we propose a framework for computing provably valid prediction bounds on claims derived from probabilistic black-box image reconstruction algorithms. The key insights are to represent reconstructed scans with a derived clinical metric of interest and to calibrate bounds on the ground truth metric with conformal prediction (CP) using a prior calibration dataset. These bounds convey interpretable feedback about the subject’s state, and can also be used to retrieve nearest-neighbor reconstructed scans for visual inspection. We demonstrate the utility of this framework on sparse-view computed tomography (CT) for fat mass quantification and radiotherapy planning tasks. Results show that our framework produces visual bounds with better semantical interpretation than conventional pixel-based bounding approaches and captures important spatial correlations. Furthermore, we can flag dangerous outlier reconstructions that look plausible but have statistically unlikely metric values. Code available at: https://github.com/matthewyccheung/conformal-metric
@@ 1 Introduction
  $s_{i,lo},s_{i,hi}]\leftarrow[Q_{\alpha_{lo}}(\hat{Y}_{i})-Y_{i},Y_{i}-Q_{1-\alpha_{hi}}(\hat{Y}_{i})]$
@@
$q_{lo},q_{hi}]\leftarrow[Q_{1-\hat{\alpha}_{lo}}(\{s_{i,lo}\}_{i=1}^{n}),Q_{1-\hat{\alpha}_{hi}}(\{s_{i,hi}\}_{i=1}^{n})]$
@@
Most existing strategies applying CP to image reconstruction operate directly on pixel values [2, 16, 29], which yield bounds with limited interpretability and do not capture spatial correlations. Other studies propose using low-dimensional representations of pixels computed using principal component analysis (PCA) [3] or latent spaces of pretrained generative models [27]. However, PCA is expensive to compute for scans larger than a small size, and high-quality generative models are not available for arbitrary medical image sets. In contrast, we propose computing CP bounds over metrics derived from the reconstructed scans, resulting in a flexible, interpretable, and computationally tractable solution.
@@ 2 Methods
For test subject $n+1$, we derive a prediction set $C(\hat{Y}_{n+1})=[L(\hat{Y}_{n+1}),U(\hat{Y}_{n+1})]$ that yields marginal coverage: $P(Y_{n+1}\in C(\hat{Y}_{n+1}))\geq 1-\alpha_{lo}-\alpha_{hi}$ [11]. If the reconstruction pipeline were perfectly calibrated, we could simply set $L(\hat{Y}_{n+1})=Q_{\alpha_{lo}}(\hat{Y}_{n+1})$ and $U(\hat{Y}_{n+1})=Q_{\alpha_{hi}}(\hat{Y}_{n+1})$. However,to handle mis-calibrations, we use $\mathcal{D}$ to find appropriate adjustments to $Q_{\alpha_{lo}}(\hat{Y}_{n+1})$ and $Q_{\alpha_{hi}}(\hat{Y}_{n+1})$. We find these adjustments in Sec. 2.1, and the use of the metric bounds to retrieve inlier, outlier, and nearest neighbor reconstructed scans in Sec. 2.2.
@@ 2.1 Quantile Adjustments Using Calibration Data
Using $\mathcal{D}$, we derive adjustments $q_{lo}$ and $q_{hi}$ to the $Q_{\alpha_{lo}}(\hat{Y}_{n+1})$ and $Q_{1-\alpha_{hi}}(\hat{Y}_{n+1})$ quantiles of the ground truth metric of test subject $n+1$ (Alg. 1). To do so, we use a probabilistic “split” CP setup proposed in previous work [32]. We first calculate $s_{lo,i}$ and $s_{hi,i}$, the lower and upper non-conformity scores for each patient, capturing the differences between the lower and upper sample quantiles and ground truth metrics: $s_{i,lo}=Q_{\alpha_{lo}}(\hat{Y}_{i})-Y_{i}$ and $s_{i,hi}=Y_{i}-Q_{1-\alpha_{hi}}(\hat{Y}_{i})$. We then compute the $(1-\alpha_{lo})$-th and $(1-\alpha_{hi})$-th empirical quantiles of these scores to arrive at the adjustments: $q_{lo}=Q_{1-\hat{\alpha}_{lo}}(\{s_{i,lo}\}_{i=1}^{n})$ and $q_{hi}=Q_{1-\hat{\alpha}_{hi}}(\{s_{i,hi}\}_{i=1}^{n})$, where $\hat{\alpha}_{lo}=\lfloor\alpha_{lo}(n+1)\rfloor/n$ and $\hat{\alpha}_{hi}=\lfloor\alpha_{hi}(n+1)\rfloor/n$ are finite-sample adjusted mis-coverage rates.
@@ 2.2 Retrieving Scans Using Metric Bounds
While the metric bounds are useful in their own right to offer interpretable feedback on patient state, they can also be mapped back to the reconstructed scans $\{\hat{Y}_{n+1,j}\}_{j=1}^{n_{s}}$ to provide visual feedback to the user or clinician (Alg. 2). First, we can flag “outlier” scans, i.e., those scans with metrics outside $L(\hat{I}_{n+1}),U(\hat{I}_{n+1})]$. Second, we can retrieve the scans with metrics nearest to $L(\hat{I}_{n+1})$ and $U(\hat{I}_{n+1})$, thereby providing visual interpretability to the metric bounds.
@@ 3 Experiments
Fat Mass Estimation. Quantifying thoracic fat mass distribution is crucial to understanding its role in cardiometabolic risk and other health outcomes [7, 31]. We computed fat volume for a 2D CT slice by multiplying the number of fat pixels with slice thickness. We identified fat pixels by thresholding values in [-150,-50] Hu. We used a de-identified CT dataset of 935 subjects from The University of Texas M.D. Anderson Cancer Center.  [^1] This research was conducted using an approved institutional review board protocol.. We split this dataset into 729 subjects ($\sim$300k axial slices) for training and 265 subjects (795 axial slices) for testing. We trained an unconditional 2D U-Net diffusion model on training data and used guided posterior sampling to reconstruct slices from 15 1-D observed projections [8, 14]. We used $n=596$ slices for calibration and 199 slices for testing, and generated $n_{s}=100$ reconstructions per slice using random input noise initializations. We repeated the experiment 1000 times with random calibration-testing splits.
@@
Table 1: Quantitative comparison of Metric CP to baselines. We used 1000 random $75\%$-$25\%$ calibration-testing splits, and report mean coverage and normalized interval lengths. The normalized interval length for a baseline method is its mean interval length divided by Metric CP’s mean interval length. While Metric, Pixel, and Pixel CP bounds often have smaller (“tighter”) interval lengths than Metric CP (a normalized length <1<1), they suffer from significant under-coverage, making them unusable for interpretable confidence assessments. By construction, Metric CP provides a target coverage with reasonable interval lengths.
@@ 4 Conclusion
We proposed a theoretically grounded, application-tuned framework to identify the quality of algorithmically reconstructed scans, and derive valid prediction intervals on derived metrics. Results show that representing scans with metrics, as opposed to operating in the raw pixel space, helps produce interpretable confidence intervals that capture meaningful spatial correlations between pixels [19, 1, 6]. However, naively computing confidence intervals directly from these metrics without calibration is insufficient, as they do not accurately reflect the true distribution of ground truth values. The proposed framework could prove useful in flagging situations where a scan-derived metric is statistically likely to be outside of acceptable safety bounds. For example, maximum dose to the heart in RT planning should be <<45Gy. If bounds for a subject indicate a good chance of the metric going beyond this limit, the case can be flagged for further review and intervention [21, 25, 33].
Our framework has a few assumptions and limitations to consider in practice. First, CP assumes exchangeability of data samples. Practitioners should ensure that calibration data is representative of future test data. Furthermore, the current method focuses on computing bounds on one metric, but there are often several metrics of interest in any application. A natural extension of this work would be to derive multi-metric bounds for image reconstruction. Moreover, image bound retrieval is only useful if there are reconstructions in the scan set with metric values sufficiently close to the metric bounds. If, for example, the reconstruction algorithm has significant bias or the number of samples $n_{s}$ is small, this may not hold and lead to misleading visual interpretations. Finally, more case studies are warranted to demonstrate the utility of these visual bounds.
{credits}
4.0.1 Acknowledgements
@@
[…]
The authors have no competing interests to declare that are relevant to the content of this article.